Harvard-Smithsonian Center for Astrophysics
FLWO Ridge Telescopes
Issue No. 1, May 2003
With this first issue, we inaugurate quarterly reports from the
F. L. Whipple Observatory. We report on the status and progress
of telescopes, instruments, related equipment and activities. The
telescopes involved are the Tillinghast 1.5m, the 1.2m and the
ex-2MASS telescope. In this issue, we also present a brief description
of the HAT (Hungarian Automated Telescope) project. This initial
report covers the period Jan 2002-Apr 2003.
All issues of these reports are posted at
Currently, the staff is led by Emilio Falco. The day-time staff are
Ted Groner (computing), Bob Hutchins (mechanical) and Wayne Peters
(electronics). Night-time staff (also known as remote observers) are
Perry Berlind and Mike Calkins. Although each member of the staff has
an assigned area of expertise, all pitch in as required in ongoing
projects, in resolution of problems as they occur, and in instrument
We concentrate here on the broader view, leaving out the day-to-day
activities required to maintain the telescopes and instruments in good
working condition. At the risk of taxing the reader's patience, this
report is comprehensive, but not necessarily perfectly inclusive. If
you have any comments or for further information on any of the
projects below, please contact us.
The instruments currently available are the 4shooter 2x2 CCD mosaic
(dark time) and the stelircam IR camera (bright time).
- We now have an FTS chiller system for the primary, with glycol
lines and two fans and heat exchangers mounted on the cell. We have
installed sensors that monitor the temperatures of the primary mirror
at several locations, as well as external and dome ambient
temperatures. The sensors are connected to an Omega control system
(fiber-isolated from the computer network and on a UPS) that will
eventually allow us to actively control the mirror temperature via
software. We display the temperatures on an interface Linux PC, based
on the system we use for stelircam. TG developed and installed a
system to view these temperatures on the web, at a link on the 1.2m
site. We are currently testing the chiller system and developing the
software to control it.
- We now have a "dog house" enclosure for the corner of the chamber
where the major heat sources (mirror chiller control and helium
compressor, hexapod control PC) reside. FLWO support staff built the
enclosure shortly after the 2002 shutdown.
- We have safety brakes installed on the telescope. After a few
mishaps such as leaky air hoses, the system is ready except for limit
switches and enable/disable controls.
- We re-aluminized the 1.2m primary and secondary in August 2002 at
Steward (we alternate yearly; the 1.5m primary was re-aluminized in
August 2001). Unfortunately, the quality of the aluminization of the
primary seems to be worsening with the mirror's age. Immediately after
the aluminization, there was a gray hazy appearance to the surface in
a few spots. The area near the center was worse than after the
previous aluminization, and the pattern is appearing in other areas of
the mirror. Steward is aware of this degradation. They minimize the
amount of time the stripping and cleaning chemicals are on the surface
of the mirror, but these still seem to be having a negative impact on
the glass. Gary Rosenbaum (Steward), who runs the aluminization
process, proposes to lightly scrub the old coating off with
hydrochloric acid instead of letting the acid soak the
coating. Monitoring the reflectivity has been hampered because the old
reflectometer is impossible to use with the mirror in the cell, and
the new Minolta reflectometer has failed.
- In April 2003 we started a program to monitor the seeing
systematically. A script is in place that allows users to automatically
measure and set the telescope focus, as well as estimate the seeing.
- Since its installation during the 2001 shutdown, the hexapod
support for the secondary mirror has performed very well. It maintains
the collimation much more robustly than the pod it replaced, and it is
a simple matter to adjust it. We monitor its performance regularly.
- Over the past two years, remote observing has become popular using
both instruments. The system has been tuned to allow that; the
internet bandwidth available has proved sufficient. The majority of
the observers have been at their Cambridge office, but we have had
cases of observers who have broadband running the telescope from the
comfort of their home.
- The infrared camera has been functioning, albeit with electronics
and temperature-related problems. The blue channel remains far from
perfect. At times it has required repeated power cycling on cold
nights. In spite of that, observers report that stelircam has produced
- We now have a system to display stelircam temperatures in graphical
form based on gnuplot. Continually-updated plots show the temperatures
of the detectors, electronics, cold head and cold structure at a
glance, which helps greatly, e.g., during warm periods that require
more frequent LN2 refills or when the He compressor has troubles.
- Because there were periodic problems, in March 2003 WP and EF
replaced the cold head with a refurbished one provided by Andy
Szentgyorgyi. We now appear to have a problem with He leaks in the
compressor system, which are still being tracked down as of this
- 4shooter recovered from a warmup that occurred in May 2002, because
its ion pump had failed. With a new pump, 4shooter returned to
normal. The new pump is not perfect; we are purchasing new ion pumps
to have enough spares for all the instruments.
- After the 2002 shutdown, we noticed a significant degradation in
4shooter flats. WP baked out the CCDs in November 2002, using a recipe
provided by John Geary. The result was much improved flats with
reduced structure, but there is secular degradation.
- Under the leadership of AS, we have proposed to upgrade the
4shooter to a monolithic 4k x 4k Fairchild CCD that will not require
UV flooding. The format is approximately equal to that of the
4shooter, which will allow recycling the mechanical and optical
designs of the 4shooter. SAO will build a new CCD controller, based on
those for the Hecto instruments and Megacam, and at least twice as
fast in readout as current controllers. An up-to-date CCD, with high
QE, without gaps and with appropriate CTE will be a great help to many
programs, in particular allowing observers to use the 1.2m to produce
astrometric data for the Hecto instruments at the MMT.
Tillinghast 1.5-meter Telescope
The instruments currently available are FAST (dark time), the Echelle
and AFOE (both bright time).
- The LKH1000 image integrator, which was used to guide with FAST
and the Echelle, started failing in early 2002. It is a model
that is no longer made or supported. The LKH1000 now resides at
ORO, where it is in use to support observations, e.g., of a pulsar.
- We replaced the LKH1000 system with a new guider system for FAST
and the Echelle, based on a frame grabber and a Linux PC. TG wrote
software for the system before and during the 2002 shutdown, including
the integration with the Telescope Control System (e.g., rotator) on
- In April 2003 we began a program to monitor the seeing
systematically with FAST. A script is in place that allows users to
automatically measure and set the telescope focus, as well as estimate
- In June 2002, we had a FAST event: novice observers left
fingerprints on the 600 gpm grating. We contacted the manufacturer,
RGL, who suggested two cleaning procedures, CO2 snow or gentle daubing
with dishwasher detergent. We attempted the CO2 first and found little
visible improvement. We decided not to attempt anything else that
would require contact with the grating and which could damage its
- In September 2002, Susan Tokarz and Mike Kurtz analyzed FAST
spectra of BG Gem for the preceding 4 years. They concluded that there
had been a significant drop in FAST sensitivity since Spring 2000 (50%
at blue wavelengths). Additional analysis of flats and standards by
PB and EF concluded that the sensitivity had indeed diminished, but by
a somewhat smaller factor. EF and RH removed and cleaned very
carefully the FAST corrector. WP baked out FAST in November 2002,
following the success we had with 4shooter. There was some gain, but
the sensitivity was still about 20% below the expected value. The one
significant improvement was that a large "divot" (a low-sensitivity
area clearly visiible in all the flats) was gone. We had Mike Lesser
clean the chip and check for vacuum leaks. There was no further
improvement in the sensitivity. The current plan is to replace the
FAST CCD. An SAO RE proposal was submitted for a replacement
chip. However, John Geary still had the old FAST2 chip. As of this
writing, he is working on FAST2 in Cambridge. We expect to have FAST2
at FLWO during the Spring 2003 trimester, and new electronics by the
end of the 2003 shutdown. Further improvements in 2003 are expected
from the work of AS on a new corrector, and of Warren Brown on all new
optics for the spectrograph.
- AFOE lives in a new enclosure, which was installed by Pete Nisenson
and Sylvain Korzennik.
- AFOE had its share of problems. Data since the repair of the
optical fibers (torn during a telescope runaway event) in January 2002
were of good quality. However, the instrument had intermittent CCD
problems. At the start of the Fall 2002 trimester, the CCD failed. PN
and SK were able to acquire a new chip in December 2002. As of this
writing, JG is working on the new CCD in the old AFOE Dewar.
- The venerable Echelle is still producing useful data, despite a
slew of small but pesky problems that plagued it over the past year.
For example, a grating-tilt motor failed, methanol leaked during
observations and the chiller lost significant chilling capacity. The
Echelle was repaired on site and in Cambridge. It is operating
normally as of this writing. The chiller may now be showing its age
and may require replacement. The addition of thicker insulation to its
glycol lines has helped to a limited degree.
2MASS 1.3-meter Telescope
- Because there was sufficient interest in trying to use the idle
telescope (which now belongs in full to SAO), in February 2002, EF
obtained CCD images (with an amateur-grade camera borrowed from the
Gamma Ray Project) of defocused stars. Brian McLeod analyzed the
images with his wavefront analysis program from Roddier and found that
the telescope optics are capable of producing 0.3-0.4 arcsec images.
- Following the tests, Nelson Caldwell and AS produced a design for a
new secondary and support system to allow us to use stelircam on the
telescope without modifying the camera. A proposal for IR&D funds was
successful. However, Harvard Fellow Josh Bloom became interested in
automating the telescope, mainly for GRB work, especially to use in
conjunction with SWIFT observations. An automatic system is necessary
to allow operations without increasing staffing. JB obtained funding
from Harvard to proceed with that part of the project. After a
modicum of uncertainty over funding for the telescope refurbishment,
JB obtained agreement from Mike Skrutskie to return the 2MASS camera
to FLWO, with no strings attached, thus eliminating the need to modify
the telescope. We hope to have the telescope and camera back on the
sky by the end of 2003.
- In March 2003, as part of his plan to automate the telescope
system, JB acquired a new weather station. Roger Harris and WP mounted
it, and JB installed monitoring software that displays wind direction
and speed, relative humidity, rainfall amounts, temperature and dew
point, and barometric pressure ina web page. It will benefit all the
ridge telescopes. Its results are available continuously on the web,
with links from each telescope web page and from the overall FLWO site.
- In 2002, we coordinated with Dan McKenna and Dave Harvey (Steward)
the installation of the SCIDAR instrument to measure seeing as a
function of altitude above Mt. Hopkins. They acquired data in June
2002 ("The LBT Scidar Facility: Recent Results", McKenna et al. 2002).
- With Steve Criswell, we submitted a request for an expansion of the
control room that would double its size and enlarge its access door,
thus making it fit for humans. M3 Engineering produced a
design. Construction awaits funding from SI.
- Gaspar Bakos, a Smithsonian predoctoral student, recently installed
the first HAT (Hungarian Automated Telescope), a small autonomous
observatory designed for robotic observations with minimal human
intervention. FLWO support and Ridge staff have assisted with items
ranging from pouring of concrete piers to running optical fibers for
- The site is next to and reuses the old 11m gamma-ray telescope
control room, near the 1.5m. The telescope uses a telephoto lens and a
2048x2048 pixel CCD, yielding a 9x9 degree field of view. Typically
with 2x4 minute exposure times, it records the brightnesses of 20000
stars with a few percent accuracy in the I-band magnitude range
6-13mag. The throughput is about one million individual photometric
measurement per night, or 2GB of raw data (including calibration
frames). The first HAT results are very promising.
- GB expects to install a total of 3 HATs. The two additional
telescopes were delivered to FLWO in April 2003; they await
installation at the end of the month. For more information, please
visit Gaspar's web site.
- Ridge staff assisted with SAO instrument installation and
maintenance, as required. RH spent the largest fraction of time
assisting with these activities, as his Ridge activities allowed.
- Ridge remote observers have assisted with observations for several
observers who had conflicts. That has allowed them to obtain useful
experience with the upgraded MMT and its new instruments. We plan to
continue such assistance, as their Ridge duties permit.
- We update online documentation continually as changes occur and,
often, thanks to comments from observers. Two examples are the
"Change Log", where significant updates to the telescopes and
instruments are reported, and the weather station reports of wind
speed and direction, relative humidity, temperature and dew point,
rainfall amounts, and barometric pressure.
- The consensus among Ridge and Cambridge staff is that all of
Mt. Hopkins network traffic should pass through a firewall, otherwise
its security is compromised. Although such consensus in 2002 expanded
to include all the projects on Mt. Hopkins, MMT staff decided instead
to use router ACLs (Access Control Lists) at CCIT (Center for
Computing and Information Technology, at U of A). FLWO purchased 2
Linux PCs (the second one as a backup) to use as firewalls. We will
proceed with the installation of a firewall for the mountain, although
the MMT network may be excluded.
- Dan Brocious and EF worked on containing light pollution. They
produced a simplified version of the Pima County Outdoor Lighting
Code, which would be more appropriate to Santa Cruz County (SCC). They made
presentations on light pollution and its control to local civic and
business organizations. A study session with the SCC
Board of Supervisors will take place during Spring 2003, and adoption
is expected shortly thereafter.
- EF and DB worked with SCC planning & zoning officials and with the
developer for the expansion of the Tubac Golf Resort. The developer
expressed keen interest in maintaining dark skies, in utilizing
appropriate lighting for new housing (about 240 new units) and proper
lighting, e.g., for tennis courts. We provided the proposed outdoor
lighting code and examples of proper luminaires and lights for sports.
- EF and DB began considering the Sopori Ranch development, which
will have an impact similar to that of the Canoa Ranch. Sopori Ranch
contains 13000 private acres and 46000 leased (State and Federal)
acres, 9000 of which are in Santa Cruz County and currently planned
- The active fire season in Spring-Summer 2002 prompted a review of
fire safety procedures on Mt. Hopkins at that time, with Karen Myres
and the Forest Service. We have established a checklist of steps to
protect Ridge equipment in case of a fire threat. Conditions for
Spring 2003 are still very dry and prone to fires, although not as
extreme as for the past year.
- We are currently planning an upgrade of the basecamp microwave link
to the Summit. The current choice is a 23GHz 100 Mb/s system (the
current system has 1.5Mb/s bandwidth) that will equalize the bandwidth
for the basecamp to summit link with the summit to Tucson link.
Emilio E. Falco
670 Mt. Hopkins Road
P.O. Box 97
Amado, AZ 85645 USA